OpenStudy (faman39):

In one week, Kean ate a total of 8 ounces of grains at breakfast, 14 ounces of grains at lunch, and 16 ounces of grains at dinner. Which of the following shows a simplified ratio of his servings of grain at lunch to his total weekly servings of grain? 19/7 7 :12 7 to 19 8: 11

OpenStudy (phoenixfire):

In one week how many ounces does Kean eat?

OpenStudy (faman39):

8

OpenStudy (phoenixfire):

no... in total.

OpenStudy (faman39):

ah, 38?

OpenStudy (phoenixfire):

Good, now how much does Kean eat at lunch?

OpenStudy (faman39):

14

OpenStudy (phoenixfire):

so the ratio is of ounces at lunch to total in a week is?

OpenStudy (faman39):

I got weird disgusting long decimal :(

OpenStudy (phoenixfire):

don't put it into a calculator... just write it here.

OpenStudy (phoenixfire):

\[{lunch \over total}={? \over ?}\]

OpenStudy (faman39):

14/38

OpenStudy (phoenixfire):

correct... and now can you simplify that?

OpenStudy (faman39):

10

OpenStudy (phoenixfire):

ummm, no. \[{14 \over 38}={7 \over 19}\]If you look in your options, 7 to 19 is one of them. That should be your answer.

OpenStudy (anonymous):

17/19

OpenStudy (faman39):

but do you guys get that ?

OpenStudy (faman39):

Why i cant get it still?

OpenStudy (faman39):

:(

OpenStudy (anonymous):

there r 7 day in week 7*14/7*38=14/38=7/19

OpenStudy (faman39):

Thank you so much guys ! :) i got it now

OpenStudy (phoenixfire):

the question says he ATE 8 ounces at breakfast, 14 at lunch and 16 at dinner during a week. you don't need to multiply by 7 because it states that IN A WEEK he ATE that much. 14 at lunch 38 in total. therefore 14/38, then simplify to get 7/19

OpenStudy (anonymous):

ok

OpenStudy (faman39):

But i m still getting wrong when i tried my own to check myself :(

OpenStudy (faman39):

I m totally stupid

OpenStudy (phoenixfire):

What are you trying to do that is giving you wrong answer?

OpenStudy (faman39):

ok

OpenStudy (faman39):

give me min i will be right back

OpenStudy (aravindg):

am i late ? :P

OpenStudy (faman39):

No :D

OpenStudy (faman39):

Ok what i m doing actually which giving me wrong answer?

OpenStudy (faman39):

I dont know :( what i m doing with it

OpenStudy (faman39):

i m totally sucks

OpenStudy (phoenixfire):

Did you not understand what I said earlier?

OpenStudy (faman39):

Yeah, i think so

OpenStudy (faman39):

have to divide 14 with 38?

OpenStudy (anonymous):

well 14/38 = 7/19

OpenStudy (faman39):

so why i m not getting it ?

OpenStudy (anonymous):

ok Faman take a deeeeeep breath

OpenStudy (anonymous):

and then think

OpenStudy (faman39):

ok, thank you so much for helping so much tolerating me :( i m sorry for troubling

OpenStudy (anonymous):

the question is simple he's just asking the ratio of grain at lunch to total grains

OpenStudy (faman39):

got it now :) finally right answer

OpenStudy (anonymous):

wow u have 164 fans

OpenStudy (anonymous):

i like ur profile faman

OpenStudy (faman39):

yeah, thank you :) i love making friends here and talk to the people and study with others and i m ok with biology so far but not good at all in math :)

OpenStudy (faman39):

Just trying to learn hard to be great

OpenStudy (anonymous):

ok thanks for ur medal

OpenStudy (faman39):

I wanted give you both credits for it, this site system is not letting me to do it

OpenStudy (faman39):

it is totally sucks

OpenStudy (anonymous):

alright that's ok

OpenStudy (faman39):

You both tried really hard for me :( thats y